# White-box Multimodal Jailbreaks Against Large Vision-Language Models

### v1

@@ Abstract
Recent advancements in Large Vision-Language Models (VLMs) have underscored their superiority in various multimodal tasks. However, the adversarial robustness of VLMs has not been fully explored. Existing methods mainly assess robustness through unimodal adversarial attacks that perturb images, while assuming inherent resilience against text-based attacks. Different from existing attacks, in this work we propose a more comprehensive strategy that jointly attacks both text and image modalities to exploit a broader spectrum of vulnerability within VLMs. Specifically, we propose a dual optimization objective aimed at guiding the model to generate affirmative responses with high toxicity. Our attack method begins by optimizing an adversarial image prefix from random noise to generate diverse harmful responses in the absence of text input, thus imbuing the image with toxic semantics. Subsequently, an adversarial text suffix is integrated and co-optimized with the adversarial image prefix to maximize the probability of eliciting affirmative responses to various harmful instructions. The discovered adversarial image prefix and text suffix are collectively denoted as a Universal Master Key (UMK). When integrated into various malicious queries, UMK can circumvent the alignment defenses of VLMs and lead to the generation of objectionable content, known as jailbreaks. The experimental results demonstrate that our universal attack strategy can effectively jailbreak MiniGPT-4 with a 96% success rate, highlighting the vulnerability of VLMs and the urgent need for new alignment strategies.
@@
Disclaimer: This paper may contain content that is potentially disturbing and offensive.
1 Introduction
The recent advancements in Large Vision-Language Models (VLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023) and Google’s Flamingo (Alayrac et al., 2022) have received broad attention from academics, industry, and the general public. These developments have also garnered widespread attention from AI safety researchers, who are increasingly focused on assessing the adversarial robustness of multimodal models. While VLMs are expected to exhibit the same alignment characteristics as Large Language Models (LLMs), providing helpful responses to user queries while rejecting requests that may cause harm, the integration of additional visual modality to base language models introduces novel vulnerabilities. Several methods (Carlini et al., 2024; Bagdasaryan et al., 2023; Bailey et al., 2023; Qi et al., 2023; Shayegani et al., 2023a) have successfully manipulated VLMs to produce harmful content by launching single-modality attacks on the continuous and high-dimensional visual modality. Conversely, research on purely text-based attacks is limited, attributed to a prevailing consensus regarding their ineffectiveness in breaching the defense mechanisms of well-aligned VLMs, due to the discrete and lower-dimensional nature of text (Carlini et al., 2024).
@@ 1 Introduction
Figure 1: Example of Jailbreak Attack on MiniGPT-4 (Zhu et al., 2023). The proposed Universal Master key (UMK) helps arbitrary harmful queries bypass alignment constraints.
In contrast, this paper presents an initial effort in introducing a text-image multimodal attack strategy, aiming to uncover a wide range of intrinsic vulnerabilities within VLMs. Furthermore, we have observed that while the previous unimodal universal attack strategy (Qi et al., 2023) increases response toxicity by perturbing images in the absence of text input, it also undermines the model’s adherence to instructions. Therefore, we propose a novel dual optimization objective strategy to address this issue.
@@
Specifically, we employ white-box attacks to identify a multimodal adversarial example capable of jailbreaking VLMs, which we refer to as the Universal Master Key (UMK). The proposed UMK comprises an adversarial image prefix and an adversarial text suffix. We expect that by attaching the UMK to arbitrary malicious user queries, it will circumvent the alignment defenses of VLMs and provoke the generation of objectionable content. To discover such a UMK, we start by initializing an adversarial image prefix from random noise and optimizing it without text input to maximize the model’s probability in generating harmful content, thereby imbuing the adversarial image prefix with toxic semantics. Subsequently, we introduce an adversarial text suffix for joint optimization with the image prefix to maximize the probability of affirmative model responses. This strategy is based on the intuition that encouraging the model to produce affirmative responses may increase its tendency to engage in undesired behaviors, rather than rejecting responses. Figure 1 exemplifies our proposed method’s efficacy in attacking MiniGPT-4 (Zhu et al., 2023).
@@
  • To the best of our knowledge, we are the first to introduce text-image multimodal adversarial attack against VLMs, systematically exploiting the vulnerabilities inherent in these models.
@@
Figure 2: Overview of our multimodal attack strategy: The Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$. We first optimize $X^{p}_{adv}$ to maximize the generation probability of harmful content without text input to infuse toxic semantics. Subsequently, we concatenate the malicious user query with $X^{s}_{adv}$, and jointly optimize $X^{p}_{adv}$ and $X^{s}_{adv}$ to maximize the generation probability of affirmative responses, e.g., ‘Sure, here’s instruction for doing ******** (a bad thing)’.
2 Related Work
2.1 Large Vision-Language Models
@@
2.2 Attacks Against Multimodal Models
@@
To attack multimodal models, Greshake et al. (Greshake et al., 2023) explored the effectiveness of manually injecting deceptive text into input images. In contrast, Other studies have proposed more sophisticated image-domain adversarial attack methods. Carlini et al. (Carlini et al., 2024) fixed the beginning portion of toxic target output and optimized the input images to increase its likelihood. Bagdasaryan et al. (Bagdasaryan et al., 2023) and Bailey et al. (Bailey et al., 2023) employed a similar strategy by using teacher-forcing techniques to generate the attacker-chosen text that may not be directly related to toxic content. Another white-box attack proposed by Qi et al. (Qi et al., 2023) adopts principles similar to Bagdasaryan et al. (Bagdasaryan et al., 2023), aiming to find an universal adversarial visual input. Specifically, the attack no longer focuses on specific output sentences but tries to maximize the probability of generating derogatory output from a corpus containing 66 harmful sample sentences. This strategy is inspired by Wallace et al. (Wallace et al., 2019), who also utilized optimization algorithms based on discrete search (Ebrahimi et al., 2017) to find universal adversarial triggers in token space, increasing the probability of generating a small group of harmful sentences. Since Qi et al. (Qi et al., 2023) did not provide a name for their method, we designate it as the Visual Adversarial Jailbreak Method (VAJM) for easier reference in subsequent sections. Shayegani et al. (Shayegani et al., 2023a) proposed attacking publicly available visual encoders such as CLIP (Radford et al., 2021) used in the multimodal models, thereby eliminating the need for complete white-box access. While the aforementioned approaches have demonstrated impressive results, they mainly focus on exploring the adversarial robustness under the unimodal attacks, based on the consensus that attacking in the continuous image space is more effective than attacking in the discrete token space. However, this focus may lead to underutilization of the full attack surfaces available in multimodal models. In contrast, our proposed method explores a broader range of vulnerabilities inherent in VLMs through attacks on text-image multimodalities. Moreover, the universal adversarial attack method proposed by Qi et al. (Qi et al., 2023) has been found to compromise the model’s adherence to user instructions. In comparison, our method employs dual optimization objectives aimed at guiding the model to generate affirmative responses with high toxicity, effectively alleviating this issue.
3 Methodology
3.1 Threat Model
@@
Attack Goals. We consider single-turn conversations between a malicious user and a VLM chatbot. The attacker attempts to trigger a number of harmful behaviors by circumventing VLM’s security mechanisms, e.g., generating unethical content or dangerous instructions restricted by system prompts or RHLF alignment technique. Adversary Capabilities. We assume that the attacker has white-box access to the target VLM, a reasonable assumption given the advanced capabilities and extensive pre-training knowledge inherent in state-of-the-art open-source VLMs. These models are sufficient to provide attackers with the insights necessary to generate malicious content, such as detailed instructions for making a bomb or materials that promote gender discrimination.
3.2 Proposed Attack
3.2.1 Formalization
@@
3.2.2 Methodology Intuition
VAJM (Qi et al., 2023) generates universal adversarial images by guiding the model to produce harmful content in the absence of text input, but this approach undermines the model’s ability to faithfully follow user instructions. In the development of jailbreak attacks against Large Language Models (LLMs), it is common practice to induce the model to respond to user queries in an affirmative manner, prioritizing task completion over query rejection. In manual attacks, this often involves prompting the model to begin its response with ‘sure’. Similarly, in Greedy Coordinate Gradient (GCG) (Zou et al., 2023), the optimization objective is to induce the model to affirmatively repeat user queries. However, we also discover that affirmatively repeating user queries does not necessarily lead to the generation of harmful content. The models still face conflicting objectives between task completion and benign alignment. Inspired by VAJM (Qi et al., 2023) and GCG (Zou et al., 2023), we propose a dual optimization objective that combines the advantages of both methods to maximize the probability of the model generating affirmative responses with high toxicity. Specifically, we begin by compromising the model’s ethical alignment through the infusion of toxic semantics into the adversarial image prefix. Subsequently, we jointly optimize the adversarial image prefix and adversarial text suffix to encourage the model to generate affirmative responses, with the expectation that the model will accomplish tasks in a state conducive to producing harmful content.
3.2.3 Embedding Toxic Semantics into Adversarial Image Prefix
@@
3.2.4 Text-Image Multimodal Optimization for Maximizing Affirmative Response Probability
@@
The shared feature space between the image embeddings and the text token embeddings in VLMs facilitates concurrent updates of both $X^{p}_{adv}$ and $X^{s}_{adv}$ through a single backpropagation pass. This mechanism is analogous to the simultaneous update of distinct pixels in a single image in conventional image adversarial attack. We employ Projected Gradient Descent (PGD) (Madry et al., 2017) to update the adversarial image prefix $X^{p}_{adv}$, and Greedy Coordinate Gradient (GCG) (Zou et al., 2023) to update the adversarial text suffix $X^{s}_{adv}$. GCG is the current state-of-the-art text-based attack strategy against LLMs. It utilizes the gradients with respect to the one-hot token indicators to identify a set of promising replacement candidates at each token position, and then chooses the substitution that results in the maximum loss reduction via a forward pass.
@@
0:  VLM model $f_{\theta}$, harmful sentences corpus $S:=\{s_{i}\}_{i=1}^{m}$, goal-target pairs corpus $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, image-text attack ratio $r$, batch size $b$, step size $\alpha$, number of candidates $n$, iteration counts $N_{1},N_{2}$;
1:  Initialize $X^{p}_{adv}$ with random noise;
2:  for $k=1,...,N_{1}$ do
3:     Select $b$ samples from the corpus $S$, forming $S_{k}:=\{s^{\prime}_{i}\}_{i=1}^{b}$;
4:     $\mathcal{L}(X^{p}_{adv},S_{k})=\sum_{i=1}^{b}-\log\left(p(s^{\prime}_{i}\mid X^{p}_{adv},\,\emptyset)\right)$;
5:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},S_{k})\right)\right)$;
6:  end for
7:  Initialize $X^{s}_{adv}$ using a predefined strategy;
8:  for $j=1,...,N_{2}$ do
9:     Select $b$ samples from the corpus $D$, forming $D_{j}:=\{g^{\prime}_{i},t^{\prime}_{i}\}_{i=1}^{b}$;
10:     $\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})=\sum_{i=1}^{b}-\log\left(p(t^{\prime}_{i}\mid X^{p}_{adv},\,g^{\prime}_{i}||X^{s}_{adv})\right)$;
11:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})\right)\right)$;
12:     if $j\%r==0$ then
13:        Compute coordinate gradient $\nabla_{x^{s}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
14:        Obtain a set of candidate suffixes with single-token substitution $\{X^{s}_{adv}\}^{n}$;
15:        $X^{s}_{adv}=\underset{\{X^{s}_{adv}\}^{n}}{\operatorname{argmin}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
16:     end if
17:  end for
18:  Obtain the Universal Master Key (UMK) composed of $X^{p}_{adv}$ and $X^{s}_{adv}$;
4 Experiments and Results
Figure 3: We present two examples of failed attacks. GCG-V produces benign content after affirmatively responding to malicious user requests, while VAJM generates harmful content without strictly adhering to user instructions.
4.1 Experimental Setup
4.1.1 Datasets

### v2

@@ Abstract.
Recent advancements in Large Vision-Language Models (VLMs) have underscored their superiority in various multimodal tasks. However, the adversarial robustness of VLMs has not been fully explored. Existing methods mainly assess robustness through unimodal adversarial attacks that perturb images, while assuming inherent resilience against text-based attacks. Different from existing attacks, in this work we propose a more comprehensive strategy that jointly attacks both text and image modalities to exploit a broader spectrum of vulnerability within VLMs. Specifically, we propose a dual optimization objective aimed at guiding the model to generate highly toxic affirmative responses. Our attack method begins by optimizing an adversarial image prefix from random noise to generate diverse harmful responses in the absence of text input, thus imbuing the image with toxic semantics. Subsequently, an adversarial text suffix is integrated and co-optimized with the adversarial image prefix to maximize the probability of eliciting affirmative responses to various harmful instructions. The discovered adversarial image prefix and text suffix are collectively denoted as a Universal Master Key (UMK). When integrated into various malicious queries, UMK can circumvent the alignment defenses of VLMs and lead to the generation of objectionable content, known as jailbreaks. The experimental results demonstrate that our universal attack strategy can effectively jailbreak MiniGPT-4 with a 96% success rate, highlighting the fragility of VLMs and the exigency for new alignment strategies. Codes are available at https://github.com/roywang021/UMK.
@@
Disclaimer: This paper contains potentially disturbing and offensive content.
Multimodal Jailbreak, White-box Jailbreak, Universal Adversarial Attacks, Vision-Language Models
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681092isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Figure 1. Example of Jailbreak Attack on MiniGPT-4 (Zhu et al., 2023). The proposed Universal Master key (UMK) helps arbitrary harmful queries bypass alignment constraints.
The recent advancements in Large Vision-Language Models (VLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023) and Google’s Flamingo (Alayrac et al., 2022) have received broad attention from academics, industry, and the general public. These developments have also garnered widespread attention from AI safety researchers, who are increasingly focused on assessing the adversarial robustness of multimodal models (Schlarmann and Hein, 2023; Zhao et al., 2024). While VLMs are expected to exhibit the same alignment characteristics as Large Language Models (LLMs), providing helpful responses to user queries while rejecting requests that may cause harm, the integration of additional visual modality to base language models introduces novel vulnerabilities. Several methods (Carlini et al., 2024; Bagdasaryan et al., 2023; Bailey et al., 2023; Qi et al., 2023; Shayegani et al., 2023a) have successfully manipulated VLMs to produce unexpected content by launching single-modality attacks on the continuous and high-dimensional visual modality. Conversely, research on purely text-based attacks is limited, attributed to a prevailing consensus regarding their ineffectiveness in breaching the defense mechanisms of well-aligned VLMs, due to the discrete and lower-dimensional nature of text (Carlini et al., 2024). In contrast, this paper presents an initial effort to introduce a text-image multimodal attack strategy, aiming to uncover a wide range of intrinsic vulnerabilities within VLMs. Furthermore, we have observed that while the previous unimodal universal attack strategy (Qi et al., 2023) increases response toxicity by perturbing images in the absence of text input, it also undermines the model’s adherence to instructions. Therefore, we propose a novel dual optimization objective strategy to address this issue.
@@
Specifically, we employ white-box attacks to identify a multimodal adversarial example capable of jailbreaking VLMs, which we refer to as the Universal Master Key (UMK). The proposed UMK comprises an adversarial image prefix and an adversarial text suffix. We expect that by attaching the UMK to arbitrary malicious user queries, it will circumvent the alignment defenses of VLMs and provoke the generation of objectionable content. To discover such a UMK, we start by initializing an adversarial image prefix from random noise and optimizing it without text input to maximize the model’s probability of generating harmful content, thereby imbuing the adversarial image prefix with toxic semantics. Subsequently, we introduce an adversarial text suffix for joint optimization with the image prefix to maximize the probability of affirmative model responses. This strategy is based on the intuition that encouraging the model to produce affirmative responses may increase its tendency to engage in undesired behaviors, rather than rejecting responses. Figure 1 exemplifies our proposed method’s efficacy in attacking MiniGPT-4 (Zhu et al., 2023).
@@
  • To the best of our knowledge, we are the first to introduce text-image multimodal adversarial attacks against VLMs, systematically exploiting the vulnerabilities inherent in these models.
@@
Figure 2. Overview of our multimodal attack strategy: The Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$. We first optimize $X^{p}_{adv}$ to maximize the generation probability of harmful content without text input to infuse toxic semantics. Subsequently, we concatenate the malicious user query with $X^{s}_{adv}$, and jointly optimize $X^{p}_{adv}$ and $X^{s}_{adv}$ to maximize the generation probability of affirmative responses, e.g., ‘Sure, here’s instruction for doing ******** (a bad thing)’.
2. Related Work
2.1. Large Vision-Language Models
@@
2.2. Attacks Against Multimodal Models
@@
To attack multimodal models, Greshake et al. (Greshake et al., 2023) explored the effectiveness of manually injecting deceptive text into input images. Gong et al. (Gong et al., 2023) converted the harmful content into images through typography to bypass the safety alignment. In contrast, Other studies have proposed more sophisticated image-domain adversarial attack methods. Carlini et al. (Carlini et al., 2024) fixed the beginning portion of toxic target output and optimized the input images to increase its likelihood. Bagdasaryan et al. (Bagdasaryan et al., 2023) and Bailey et al. (Bailey et al., 2023) employed a similar strategy by using teacher-forcing techniques to generate the attacker-chosen text that may not be directly related to toxic content. Another white-box attack proposed by Qi et al. (Qi et al., 2023) adopts principles similar to Bagdasaryan et al. (Bagdasaryan et al., 2023), aiming to find a universal adversarial visual input. Specifically, the attack no longer focuses on specific output sentences but tries to maximize the probability of generating derogatory output from a corpus containing 66 harmful sample sentences. This strategy is inspired by Wallace et al. (Wallace et al., 2019), who also utilized optimization algorithms based on discrete search (Ebrahimi et al., 2017) to find universal adversarial triggers in token space, increasing the probability of generating a small group of harmful sentences. Since Qi et al. (Qi et al., 2023) did not provide a name for their method, we designate it as the Visual Adversarial Jailbreak Method (VAJM) for easier reference in subsequent sections. Concurrent and independent from our work, Niu et al. (Niu et al., 2024) adopted the same optimization objective as our second stage to generate adversarial images on VLMs, then endeavored to transform them into adversarial text suffixes for LLMs. While the aforementioned approaches have demonstrated impressive results, they mainly focus on exploring the adversarial robustness under the unimodal attacks. In contrast, our proposed method explores a broader range of vulnerabilities inherent in VLMs through attacks on text-image multimodalities. Additionally, the universal adversarial attack method proposed by Qi et al. (Qi et al., 2023) has been found to compromise the model’s adherence to user instructions and the single optimization objective of producing affirmative responses, as employed in concurrent work (Niu et al., 2024), could result in insufficient toxicity in the model’s output. In comparison, our method utilizes dual optimization objectives aimed at guiding the model to generate highly toxic affirmative responses, effectively alleviating this concern.
3. Methodology
3.1. Threat Model
@@
Attack Goals. We consider single-turn conversations between a malicious user and a VLM chatbot. The attacker attempts to trigger a number of harmful behaviors by circumventing VLM’s security mechanisms, e.g., generating unethical content or dangerous instructions restricted by system prompts or the RHLF alignment technique. Adversary Capabilities. We assume that the attacker has white-box access to the target VLM, a reasonable assumption given the advanced capabilities and extensive pre-training knowledge inherent in state-of-the-art open-source VLMs. These models are sufficient to provide attackers with the insights necessary to generate malicious content, such as detailed instructions for making a bomb or materials that promote gender discrimination.
3.2. Proposed Attack
3.2.1. Formalization
@@
3.2.2. Methodology Intuition
VAJM (Qi et al., 2023) generates universal adversarial images by guiding the model to produce harmful content in the absence of text input, but this approach undermines the model’s ability to faithfully follow user instructions. In the development of jailbreak attacks against Large Language Models (LLMs), it is common practice to induce the model to respond to user queries in an affirmative manner, prioritizing task completion over query rejection. In manual attacks, this often involves prompting the model to begin its response with ‘sure’. Similarly, in Greedy Coordinate Gradient (GCG) (Zou et al., 2023), the optimization objective is to induce the model to affirmatively repeat user queries. However, we also discover that affirmatively repeating user queries does not necessarily lead to the generation of harmful content. The models still face conflicting objectives between task completion and benign alignment. Inspired by VAJM (Qi et al., 2023) and GCG (Zou et al., 2023), we propose a dual optimization objective that combines the advantages of both methods to maximize the probability of the model generating highly toxic affirmative responses. Specifically, we begin by compromising the model’s ethical alignment through the infusion of toxic semantics into the adversarial image prefix. Subsequently, we jointly optimize the adversarial image prefix and adversarial text suffix to encourage the model to generate affirmative responses, with the expectation that the model will accomplish tasks in a state conducive to producing harmful content.
3.2.3. Embedding Toxic Semantics into Adversarial Image Prefix
@@
3.2.4. Text-Image Multimodal Optimization for Maximizing Affirmative Response Probability
@@
The shared feature space between the image embeddings and the text token embeddings in VLMs facilitates concurrent updates of both $X^{p}_{adv}$ and $X^{s}_{adv}$ through a single backpropagation pass. This mechanism is analogous to the simultaneous update of distinct pixels in a single image in a conventional image adversarial attack. We employ Projected Gradient Descent (PGD) (Madry et al., 2017) to update the adversarial image prefix $X^{p}_{adv}$, and Greedy Coordinate Gradient (GCG) (Zou et al., 2023) to update the adversarial text suffix $X^{s}_{adv}$. GCG is the current state-of-the-art text-based attack strategy against LLMs. It utilizes the gradients with respect to the one-hot token indicators to identify a set of promising replacement candidates at each token position and then chooses the substitution that results in the maximum loss reduction via a forward pass.
@@
1:VLM model $f_{\theta}$, harmful sentences corpus $S:=\{s_{i}\}_{i=1}^{m}$, goal-target pairs corpus $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, image-text attack ratio $r$, batch size $b$, step size $\alpha$, number of candidates $n$, iteration counts $N_{1},N_{2}$;
2:Initialize $X^{p}_{adv}$ with random noise;
3:for $k=1,...,N_{1}$ do
4:     Select $b$ samples from the corpus $S$, forming $S_{k}:=\{s^{\prime}_{i}\}_{i=1}^{b}$;
5:     $\mathcal{L}(X^{p}_{adv},S_{k})=\sum_{i=1}^{b}-\log\left(p(s^{\prime}_{i}\mid X^{p}_{adv},\,\emptyset)\right)$;
6:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},S_{k})\right)\right)$;
7:end for
8:Initialize $X^{s}_{adv}$ using a predefined strategy;
9:for $j=1,...,N_{2}$ do
10:     Select $b$ samples from the corpus $D$, forming $D_{j}:=\{g^{\prime}_{i},t^{\prime}_{i}\}_{i=1}^{b}$;
11:     $\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})=\sum_{i=1}^{b}-\log\left(p(t^{\prime}_{i}\mid X^{p}_{adv},\,g^{\prime}_{i}||X^{s}_{adv})\right)$;
12:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})\right)\right)$;
13:     if $j\%r==0$ then
14:         Compute coordinate gradient $\nabla_{x^{s}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
15:         Obtain a set of candidate suffixes with single-token substitution $\{X^{s}_{adv}\}^{n}$;
16:         $X^{s}_{adv}=\underset{\{X^{s}_{adv}\}^{n}}{\operatorname{argmin}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
17:     end if
18:end for
19:Obtain the Universal Master Key (UMK) composed of $X^{p}_{adv}$ and $X^{s}_{adv}$;
4. Experiments and Results
Figure 3. We present two examples of failed attacks. GCG-V produces benign content after affirmatively responding to malicious user requests, while VAJM generates harmful content without strictly adhering to user instructions.
4.1. Experimental Setup
4.1.1. Datasets